OpenStudy (anonymous):

Find the height of a rectangular prism if the surface area is 868, the width is 7 and the length is 31. (Points : 2) 434 2.85 76 5.7

OpenStudy (anonymous):

The height will be the surface area divided by the perimeter of the base. You have the base dimensions and the surface area, so just do the calculation.

OpenStudy (anonymous):

So, 124?

OpenStudy (anonymous):

Use: \[\large A = 2[l \cdot h + h \cdot w +w \cdot l ]\]

OpenStudy (anonymous):

But I need to know the height not area